OpenStudy (ali2x2):

2 + 2 - 7 x 0 + 4 - 32 x 89

OpenStudy (anonymous):

-2840

undeadknight26 (undeadknight26):

4 - 32 x 89 Actual equation

OpenStudy (ali2x2):

seasal

OpenStudy (ali2x2):

staph using calc

undeadknight26 (undeadknight26):

@saseal is a human calculator.

OpenStudy (anonymous):

waddup?

OpenStudy (ali2x2):

bro i can clearly see your using calc

OpenStudy (anonymous):

of course i am

OpenStudy (ali2x2):

:P

OpenStudy (ali2x2):

ok seasal you make a question

OpenStudy (ali2x2):

post a question we will go to it

OpenStudy (anonymous):

what kind of level of questions?

OpenStudy (ali2x2):

ayy stupid questions that involve x, -, +, divide

undeadknight26 (undeadknight26):

Too lazy ima come off now xD I'm going to pick up the car today

OpenStudy (ali2x2):

\(\color{#0cbb34}{\text{Originally Posted by}}\) @undeadknight26 Too lazy ima come off now xD I'm going to pick up the car today \(\color{#0cbb34}{\text{End of Quote}}\) AYYYYY

OpenStudy (anonymous):

6 x 2 + 20 + 9 * 0 +89

OpenStudy (ali2x2):

12 + 20 = 32

OpenStudy (ali2x2):

+ 9 =41

OpenStudy (anonymous):

nope

OpenStudy (ali2x2):

what?

OpenStudy (anonymous):

do the whole thing first

OpenStudy (ali2x2):

oh 31*

undeadknight26 (undeadknight26):

89

OpenStudy (ali2x2):

31 x 0 = 0

OpenStudy (ali2x2):

STAPH USING CALC

OpenStudy (anonymous):

its not 89

undeadknight26 (undeadknight26):

I didn't use calc I just guessed x.x

OpenStudy (ali2x2):

it is 89...

OpenStudy (anonymous):

its 121

undeadknight26 (undeadknight26):

6 x 2 + 20 + 9 * 0 +89 12 + 20 + 0 + 89

OpenStudy (anonymous):

6 X 2 +20 + 89 = 12 + 20 +89 = 121

OpenStudy (ali2x2):

no you didnt use ( )

OpenStudy (anonymous):

what kinda calc ya'll use?

OpenStudy (ali2x2):

the simplest

OpenStudy (ali2x2):

my mind .-.

OpenStudy (anonymous):

lel noob

OpenStudy (ali2x2):

nu i no n00b

OpenStudy (ali2x2):

1v1 m3 i am grammer naz1 i am b3st sp3ll3r i w0n the sp3ll1ng b33

OpenStudy (anonymous):

░░░░░░░░░▄░░░░░░░░░░░░░░▄░░░░ ░░░░░░░░▌▒█░░░░░░░░░░░▄▀▒▌░░░ ░░░░░░░░▌▒▒█░░░░░░░░▄▀▒▒▒▐░░░ ░░░░░░░▐▄▀▒▒▀▀▀▀▄▄▄▀▒▒▒▒▒▐░░░ ░░░░░▄▄▀▒░▒▒▒▒▒▒▒▒▒█▒▒▄█▒▐░░░ ░░░▄▀▒▒▒░░░▒▒▒░░░▒▒▒▀██▀▒▌░░░ ░░▐▒▒▒▄▄▒▒▒▒░░░▒▒▒▒▒▒▒▀▄▒▒▌░░ ░░▌░░▌█▀▒▒▒▒▒▄▀█▄▒▒▒▒▒▒▒█▒▐░░ ░▐░░░▒▒▒▒▒▒▒▒▌██▀▒▒░░░▒▒▒▀▄▌░ ░▌░▒▄██▄▒▒▒▒▒▒▒▒▒░░░░░░▒▒▒▒▌░ ▀▒▀▐▄█▄█▌▄░▀▒▒░░░░░░░░░░▒▒▒▐░ ▐▒▒▐▀▐▀▒░▄▄▒▄▒▒▒▒▒▒░▒░▒░▒▒▒▒▌ ▐▒▒▒▀▀▄▄▒▒▒▄▒▒▒▒▒▒▒▒░▒░▒░▒▒▐░ ░▌▒▒▒▒▒▒▀▀▀▒▒▒▒▒▒░▒░▒░▒░▒▒▒▌░ ░▐▒▒▒▒▒▒▒▒▒▒▒▒▒▒░▒░▒░▒▒▄▒▒▐░░ ░░▀▄▒▒▒▒▒▒▒▒▒▒▒░▒░▒░▒▄▒▒▒▒▌░░ ░░░░▀▄▒▒▒▒▒▒▒▒▒▒▄▄▄▀▒▒▒▒▄▀░░░ ░░░░░░▀▄▄▄▄▄▄▀▀▀▒▒▒▒▒▄▄▀░░░░░ ░░░░░░░░░▒▒▒▒▒▒▒▒▒▒▀▀░░░░░░░░

OpenStudy (anonymous):

my selfie